The time between arrivals of 52 patients at an intensive care unit were recorded to the nearest hour. The data are shown below. (20 Marks)




Time (Hours)



[0-20)



[20-40)



[40-60)



[60-80)



[80-100)



[100-120)



[120-140)



Frequency



16



13



7



4



3



4



5





I. Construct a histogram and cumulative frequency polygon for these data.



II. Estimate the percentage of the time between arrivals of patients with less than 80 hours.

1
Expert's answer
2022-01-04T08:03:22-0500

i.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Time(hours) & Cumulative\ frequency \\ \hline
 0 & 0 \\
 19.5 & 16\\
 39.5 & 29\\
 59.5 & 36\\
 79.5 & 40\\
 99.5 & 43\\
 119.5 & 47\\
 139.5 & 52\\
\end{array}"



ii.


"P(T<80)=\dfrac{40}{52}=\dfrac{10}{13}\approx0.76923"

"77\%"